# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

### v2

@@ Abstract.
Diffusion Models (DMs) have achieved remarkable success in realistic voice cloning (VC), while they also increase the risk of malicious misuse. Existing proactive defenses designed for traditional VC models aim to disrupt the forgery process, but they have been proven incompatible with DMs due to the intricate generative mechanisms of diffusion. To bridge this gap, we introduce VoiceCloak, a multi-dimensional proactive defense framework with the goal of obfuscating speaker identity and degrading perceptual quality in potential unauthorized VC. To achieve these goals, we conduct a focused analysis to identify specific vulnerabilities within DMs, allowing VoiceCloak to disrupt the cloning process by introducing adversarial perturbations into the reference audio. Specifically, to obfuscate speaker identity, VoiceCloak first targets speaker identity by distorting representation learning embeddings to maximize identity variation, which is guided by auditory perception principles. Additionally, VoiceCloak disrupts crucial conditional guidance processes, particularly attention context, thereby preventing the alignment of vocal characteristics that are essential for achieving convincing cloning. Then, to address the second objective, VoiceCloak introduces score magnitude amplification to actively steer the reverse trajectory away from the generation of high-quality speech. Noise-guided semantic corruption is further employed to disrupt structural speech semantics captured by DMs, degrading output quality. Extensive experiments highlight VoiceCloak’s outstanding defense success rate against unauthorized diffusion-based voice cloning.
Digital forensics, Proactive defense, Diffusion models, Adversarial examples, Voice clone
ccs: Security and privacy Social aspects of security and privacyccs: Applied computing Sound and music computingcopyright: none
1. Introduction
@@
Diffusion Models (DMs) (Ho et al., 2020; Song et al., 2020a, b; Rombach et al., 2022) have recently emerged as powerful generative tools, achieving unprecedented success within speech synthesis, particularly for realistic voice cloning (VC). Their iterative denoising process enables fine-grained control over the generation procedure, yielding synthetic speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al., 2022; Kong et al., 2020; Jeong et al., 2021; Shen et al., 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
@@
Figure 1. Illustration of diffusion-based voice cloning malicious misuse. (a) Voice forgery enables threats of fraud. (b) Traditional methods struggle due to ineffective disruptive gradients. (c) Audio protected by VoiceCloak resists high-fidelity cloning.
@@
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Existing forgery detection methods for synthesized speech (Jung et al., 2022; Zhou and Lim, 2021; Wu et al., 2024) are reactive, often failing to prevent harm after dissemination. This highlights a critical need for proactive defenses that mitigate these security risks before they manifest. To this end, several studies (Huang et al., 2021; Yu et al., 2023; Chen et al., 2024; Li et al., 2023) have explored proactive disruption by introducing elaborately crafted, imperceptible, adversarial perturbations into reference audio. These perturbations aim to compromise the functionality of either the voice decoder or the speaker identity encoder, thereby disrupting the cloning process.
@@
While effective on traditional conversion architectures, existing proactive defense methods fail to defend against diffusion-based voice cloning. Figure 1(b) illustrates that the incompatibility arises from gradient vanishing and dynamic conditioning, which present unique challenges for traditional defense methods. Specifically, (1) Established methods rely on computing gradients via a single forward pass of the generator, aligning perturbations with the generator’s complete structure. However, the DMs’ long, multi-step sampling process creates a deep computational graph prone to vanishing gradients. This makes single-pass gradient information ineffective at disrupting the full generation trajectory (Kang et al., 2023). (2) Other strategies target specific subnetworks, such as encoders designed to process inputs like identity or content independently locally. Conversely, DMs often employ dynamic conditioning mechanisms, where the focus can shift across conditional aspects. Consequently, localized methods designed to target static modules struggle to achieve global disruption in DMs. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
@@
In light of the incompatibility of prior defenses with the diffusion paradigm, we introduce VoiceCloak, a novel proactive defense framework designed to achieve two primary objectives against unauthorized Diffusion-based Voice Cloning (VC): Speaker Identity Obfuscation and Perceptual Fidelity Degradation. Driven by these goals, we conduct a targeted analysis to identify corresponding intrinsic vulnerabilities within the Diffusion Models (DMs). Based on this analysis, we design specific optimization objectives for the adversarial perturbation. The perturbation then exploits analysis results to effectively disrupt the diffusion-based voice cloning pipeline.
@@
To achieve identity obfuscation, VoiceCloak first directly manipulates speaker identity representations within a universal embedding space to prevent DMs from extracting the target identity signature. We further optimize a loss term derived from psychoacoustic principles to maximize perceptual identity distance. Recognizing that convincing mimicry requires precise conditional guidance via attention mechanisms. This dependency makes attention a key for disrupting identity rendering. To exploit this, we introduce attention context divergence to prevent the attention mechanism from correctly utilizing the contextual information derived from the reference audio, thus disrupting the accurate alignment of speaker style with content and hindering convincing cloning.
@@
Simultaneously, to achieve fidelity degradation, our analysis focused on vulnerabilities within the core generative process itself. We identified the precision of the iterative denoising trajectory as crucial for high-fidelity generation. Therefore, we employ score magnitude amplification (SMA) to exploit this sensitivity, actively steering the generation path away from high-quality speech regions. Additionally, acknowledging that the U-Net’s internal understanding of structural speech semantics contributes to naturalness, we utilize noise-guided semantic corruption to disrupt the capture of coherent structural features and promote incoherence within the noise semantic space, degrading the final audio quality. These goal-driven strategies, each originating from an analysis of vulnerabilities corresponding to our defense objectives, form the overall VoiceCloak framework. Extensive experiments confirm the superior defense efficacy of our method against diffusion-based voice cloning attacks under equivalent perturbation budgets.
@@
2. Related Work
2.1. Audio Diffusion Models
@@
Diffusion Models (Ruan et al., 2023; Kawar et al., 2023; Ruiz et al., 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al., 2021; Jeong et al., 2021) and unconditional audio generation (Kong et al., 2020; Liu et al., 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al., 2022; Choi et al., 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al., 2020b), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
2.2. Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, such as Voice Cloning (VC) or Text-to-Speech (TTS), by introducing carefully crafted, often imperceptible adversarial perturbations to the original audio.
Initial efforts by Huang et al. (Huang et al., 2021) demonstrated the feasibility of perturbing reference audio to disrupt VC, though balancing effectiveness and imperceptibility remained a challenge. Subsequent research sought to improve this balance. For instance, VoiceGuard (Li et al., 2023) focused on imperceptibility, optimizing time-domain perturbations guided by psychoacoustic models to stay below auditory thresholds. AntiFake (Yu et al., 2023) pursued similar goals, incorporating a human-in-the-loop process to iteratively refine the trade-off between perceived identity protection and audio quality. Addressing efficiency, Dong et al. (Dong et al., 2024) proposed a GAN-based generator to create perturbations more rapidly than typical iterative optimization methods.
@@ 2.2. Proactive Defense via Adversarial Examples
Despite these advancements, existing proactive strategies were largely developed for prior generative architectures (e.g., GANs). They often overlook the unique denoising processes and internal structures characteristic of Diffusion Models (DMs), which differ fundamentally from earlier paradigms. Recognizing this critical gap, we propose an effective defense specifically designed to counter the challenges posed by diffusion-based voice cloning.
3. Preliminaries
3.1. Score-based Diffusion
Unlike the original denoising diffusion model that utilizes a discrete-time diffusion process by Markov chains (Ho et al., 2020), the score-based generative models leverage stochastic differential equations (Song et al., 2020b) to define continuous-time diffusion processes. The forward process progressively adds noise to the original audio $x_{0}\sim p_{0}(x)$ to transform the complex data distribution of $p_{0}(x)$ into a simple prior Gaussian distribution. The evolution of noisy audio $x(t)$ is governed by an SDE:
(1) $d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w},$
@@ 3.1. Score-based Diffusion
where $f(\mathbf{x},t)$ is the drift coefficient, $g(t)$ is the diffusion coefficient controlling the noise level and $\mathbf{w}$ is a standard Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(x)$ can be expressed as:
(2) $d\mathbf{x}=\left[f(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}},$
@@
(3) $\theta^{*}=\arg\min_{\theta}\int_{0}^{1}\lambda_{t}\mathbb{E}_{\mathbf{x}_{t}}\left\|s_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|_{2}^{2}dt,$
@@
3.2. Adversarial Vulnerability Analysis
@@
A primary objective of VoiceCloak is speaker identity obfuscation. Convincing identity mimicry in Diffusion Models (DMs) necessitates precise conditional control, guided by fine-grained acoustic details perceived and modeled from the reference audio $x_{ref}$. This guidance critically relies on mechanisms that align the target speaker’s acoustic characteristics from $x_{ref}$ with the corresponding phonetic content from $x_{src}$. Specifically, the attention block within U-Net executes this alignment, establishing the crucial mapping between source content elements and relevant acoustic style features derived from $x_{ref}$. This content-to-style mapping creates a key vulnerability: inaccurate attention-based alignment directly compromises the successful rendering of the target speaker’s identity. Consequently, disrupting this pathway offers a strategy for identity obfuscation.
@@
Complementary to identity obfuscation, the second objective is perceptual fidelity degradation, intended to diminish the usability of potentially synthesized speech. This involves targeting core mechanisms underpinning the high quality achieved by DMs. Central to this is the iterative denoising process. Diffusion-based VC synthesizes high-fidelity audio by progressively refining a sample, starting from noise $x_{T}\sim\mathcal{N}(0,I)$, guided at each timestep $t$ towards the real data distribution. Consequently, a well-trained model effectively learns a precise reverse denoising trajectory. Faithfully executing this learned trajectory is essential for generating the natural outputs. This reliance on trajectory precision presents an exploitable vulnerability. Therefore, adversarially diverting the trajectory can disrupt convergence towards the high-fidelity audio region on the manifold.
@@
Additionally, the perceived quality and naturalness of synthesized speech are closely tied to the coherent processing of structural information within the U-Net architecture (Ronneberger et al., 2015). The U-Net generates hierarchical internal feature maps capturing essential data patterns. Prior attribute editing work (Oh et al., 2024; Tumanyan et al., 2023) confirms these internal features are controllable and encode crucial semantic and acoustic details impacting perceptual quality. Adversarially corrupting these representations can directly impair the model’s capacity to synthesize coherent and natural-sounding speech. Therefore, targeting these U-Net features offers a complementary strategy. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
4. Methodology
@@
Figure 2. Overview of the proposed VoiceCloak framework. Perturbation optimization is guided by gradients from $\mathcal{L}_{total}$, aggregating contributions from four core components targeting two defense objective: (1) Identity Obfuscation, achieved via Opposite-Gender Centroid Guidance and Attention Context Divergence; and (2) Perceptual Fidelity Degradation, implemented through Score Magnitude Amplification and Noise-Guided Semantic Corruption.
@@
In this section, we elaborate on the design and implementation details of the proposed VoiceCloak framework. To provide the necessary background, the problem formulation is established in Section 4.1. As shown in Figure 2, the VoiceCloak consists of four sub-modules, which will be introduced separately next.
4.1. Problem Formulation
@@
Assume that malicious users can obtain reference audio $x_{ref}$ of a target speaker from public platforms. Leveraging open-source Voice Cloning (VC) models, they can synthesize speech from $x_{src}$ that mimics the vocal characteristics of $x_{ref}$, denoted $\mathcal{VC}(x_{src},x_{ref},t)$. Our goal is to proactively safeguard $x_{ref}$ against unauthorized voice cloning. The core idea is to introduce a human-imperceptible perturbation $\delta$ into $x_{ref}$, generating an adversarial audio $x_{adv}=x_{ref}+\delta$, which disrupts the reverse diffusion process leading to corrupted conversions. Formally, our objective is to find an optimal perturbation $\delta$ that maximizes the dissimilarity between the outputs generated using $x_{ref}$ and $x_{adv}$:
(4) $\delta:=$ $\arg\max_{\delta}\ \mathcal{D}(\mathcal{VC}(x_{src},x_{ref},t),\ \mathcal{VC}(x_{src},x_{adv},t)),$
subject $\text{subject to}\quad\|\delta\|_{\infty}\leq\epsilon,$
@@
where $\mathcal{D}(\cdot)$ is a function that measures the discrepancy, $t$ indicates the diffusion timestep, $\epsilon$ is the infinity norm bound budget to restrict $\delta$. Notably, we select DiffVC (Popov et al., 2022) as our primary threat model due to its representativeness in diffusion-based voice cloning. Our approach focuses on designing specific adversarial objectives that implicitly define $\mathcal{D}$ by exploiting the intrinsic vulnerabilities.
4.2. Adversarial Identity Obfuscation
Opposite-Gender Embedding Centroid Guidance.

### v4

@@ Abstract
Diffusion Models (DMs) have achieved remarkable success in realistic voice cloning (VC), while they also increase the risk of malicious misuse. Existing proactive defenses designed for traditional VC models aim to disrupt the forgery process, but they have been proven incompatible with DMs due to the intricate generative mechanisms of diffusion. To bridge this gap, we introduce VoiceCloak, a multi-dimensional proactive defense framework with the goal of obfuscating speaker identity and degrading perceptual quality in potential unauthorized VC. To achieve these goals, we conduct a focused analysis to identify specific vulnerabilities within DMs, allowing VoiceCloak to disrupt the cloning process by introducing adversarial perturbations into the reference audio. Specifically, to obfuscate speaker identity, VoiceCloak first targets speaker identity by distorting representation learning embeddings to maximize identity variation, which is guided by auditory perception principles. Additionally, VoiceCloak disrupts crucial conditional guidance processes, particularly attention context, thereby preventing the alignment of vocal characteristics that are essential for achieving convincing cloning. Then, to address the second objective, VoiceCloak introduces score magnitude amplification to actively steer the reverse trajectory away from the generation of high-quality speech. Noise-guided semantic corruption is further employed to disrupt structural speech semantics captured by DMs, degrading output quality. Extensive experiments highlight VoiceCloak’s outstanding defense success rate against unauthorized diffusion-based voice cloning. Audio samples of VoiceCloak are available at https://voice-cloak.github.io/VoiceCloak/.
1 Introduction
@@
Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Rombach et al. 2022) have recently emerged as powerful generative tools, achieving unprecedented success within realistic voice cloning (VC). Their iterative denoising process enables generating speech with remarkable naturalness, detail, and fidelity to human voice (Popov et al. 2022; Kong et al. 2020; Jeong et al. 2021; Shen et al. 2024). However, the open-source availability and ease of use of these models intensify concerns about potential misuse. Attackers can synthesize highly realistic voice replicas from short public audio clips, as depicted in Figure 1(a), enabling sophisticated fraud and circumvention of voiceprint authentication.
@@
Figure 1: Illustration of diffusion-based voice cloning malicious misuse. (a) Voice forgery enables threats of fraud. (b) Traditional methods struggle due to ineffective disruptive gradients. (c) Audio protected by VoiceCloak resists high-fidelity cloning.
@@
To counter such unauthorized use, two main defense paradigms are introduced, including forgery detection and proactive disruption. Reactive detection methods (Jung et al. 2022; Zhou and Lim 2021; Wu et al. 2024) identify forgeries after they are crafted, often too late to prevent harm. This highlights the need for proactive defenses that disrupt the synthesis process itself. Prior proactive work (Huang et al. 2021; Yu, Zhai, and Zhang 2023; Chen et al. 2024; Li et al. 2023) has focused on adding imperceptible adversarial perturbations to reference audio by compromising the functionality of either the voice decoder or the speaker identity encoder.
@@
However, existing defenses designed for prior architectures are largely ineffective against Diffusion Models (DMs). This incompatibility arises from two fundamental challenges: gradient vanishing and dynamic conditioning (Figure 1). Specifically, (1) the single forward pass gradient computation relied upon by many defenses become unreliable or vanish within the multi-step denoising process of DMs and the corresponding deep computational graph, rendering such single-pass gradient information ineffective for disrupting the full generation trajectory (Kang, Song, and Li 2023). (2) Strategies targeting specific subnetworks (e.g. speaker or content encoders) fail because DMs often employ dynamic conditioning mechanisms, which means no single modules solely responsible for condition processing. Consequently, methods targeting individual components struggle to cause global disruption. These fundamental incompatibilities underscore the need for novel strategies tailored to this generative paradigm.
@@
In light of the incompatibility of prior defenses with the diffusion paradigm, we introduce VoiceCloak, a novel proactive defense framework designed for two primary objectives against unauthorized voice cloning: Speaker Identity Obfuscation and Perceptual Fidelity Degradation. Driven by these two objectives, we conduct an analysis to identify and exploit corresponding intrinsic vulnerabilities within Diffusion Models (DMs). Based on this analysis, we design specific optimization objectives for the protective perturbation to effectively disrupt the synthesis process.
@@
To achieve identity obfuscation, VoiceCloak first directly manipulates speaker representations within a universal embedding space, guided by psychoacoustic principles to maximize perceived identity distance and hinder the DM’s identity signature extraction. Second, recognizing that convincing mimicry depends on attention mechanisms in conditional guidance to align speaker style with content, we exploit this by introducing attention context divergence. This design prevents the attention mechanism from correctly utilizing contextual information, thereby disrupting the alignment required for accurate cloning.
@@
Simultaneously, to achieve fidelity degradation, we focus on vulnerabilities within the core generative process itself. First, we employ Score Magnitude Amplification (SMA) to exploit the sensitivity of the iterative denoising trajectory which is crucial for realistic output, steering the generation path away from high-fidelity regions. Furthermore, acknowledging that the U-Net’s understanding of high-level semantics governs output naturalness, we utilize noise-guided semantic corruption to disrupt the capture of structural features to promote incoherence within the noise semantic space and degrade generation quality. These goal-driven strategies which originates from adversarial analysis form a comprehensive defense. Extensive experiments confirm the superior defense efficacy against diffusion-based VC attacks under equivalent perturbation budgets.
@@
2 Related Work
2.1 Audio Diffusion Models
@@
Diffusion Models (Ruan et al. 2023; Kawar et al. 2023; Ruiz et al. 2023) a dominant force in generative modeling, demonstrating extraordinary performance across multi-modal tasks, significantly advancing audio synthesis tasks like text-to-speech (Popov et al. 2021; Jeong et al. 2021) and unconditional audio generation (Kong et al. 2020; Liu et al. 2023). Particularly within voice cloning (VC), diffusion-based methods(Popov et al. 2022; Choi, Lee, and Lee 2024), mostly leveraging score-based formulations via stochastic differential equations (SDEs) (Song et al. 2020), now yield outputs with remarkable naturalness and speaker fidelity. While impressive, this state-of-the-art performance significantly heightens concerns regarding potential misuse, directly motivating proactive defense strategies such as the framework proposed herein.
2.2 Proactive Defense via Adversarial Examples
Beyond passive DeepFake detection, proactive defenses aim to preemptively disrupt malicious syntheses, by introducing adversarial perturbations to the original audio. Early work (Huang et al. 2021) demonstrated the feasibility of this approach but struggled to balance effectiveness with imperceptibility. Subsequent research focused on improving this trade-off. Strategies included using psychoacoustic models to enhance imperceptibility (Li et al. 2023), incorporating human-in-the-loop refinement for better balance (Yu, Zhai, and Zhang 2023), , and employing GAN-based generators to improve efficiency (Dong et al. 2024).
@@ 2.2 Proactive Defense via Adversarial Examples
Despite these advancements, prior proactive strategies were largely designed for earlier generative architectures like GANs. They often overlook the unique mechanisms and internal structures of Diffusion Models (DMs), limiting their applicability. Recognizing this critical gap, our work proposes a defense specifically tailored to the challenges of diffusion-based voice cloning.
3 Preliminaries
3.1 Score-based Diffusion
Score-based generative models define a continuous-time diffusion process using stochastic differential equations (SDEs) (Song et al. 2020). The forward process gradually perturbs clean data $x_{0}\sim p_{0}(x)$ into noise via the SDE:
$d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w},$ (1)
@@ 3.1 Score-based Diffusion
where $f(\mathbf{x},t)$ and $g(t)$ are the drift and diffusion coefficients, and $\mathbf{w}$ is a Wiener process. The corresponding reverse-time SDE that transforms $x_{T}$ back into $p_{0}(x)$ can be expressed as:
$d\mathbf{x}=\left[f(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}},$ (2)
@@
$\arg\min_{\theta}\lambda_{t}\mathbb{E}_{p_{t}(x)}\left\|s_{\theta}(\mathbf{x}_{t},t)-\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}_{2},$ (3)
@@
3.2 Adversarial Vulnerability Analysis
@@
A primary objective of VoiceCloak is speaker identity obfuscation. Convincing identity mimicry in Diffusion Models (DMs) depends on precise conditional control which is guided by acoustic details modeled from the reference audio $x_{ref}$. This guidance critically relies on mechanisms that align the target speaker’s acoustic characteristics from $x_{ref}$ with the phonetic content from $x_{src}$. Specifically, the attention block is responsible for executing this alignment. We identify this crucial acoustic style-to-content mapping as a key vulnerability,as inaccurate alignment directly compromises the successful rendering of the target speaker’s identity. Consequently, disrupting this attention-driven pathway offers a strategy for identity obfuscation.
@@
Complementary to identity obfuscation, the second objective is perceptual fidelity degradation, aimed at diminishing the usability of any synthesized speech. This involves targeting the core denoising process of DMs. The high fidelity of Diffusion-based VC relies on the model learning a precise reverse denoising trajectory to progressively refine noisy sample $x_{t}$ towards the natural audio distribution. This reliance on trajectory precision presents an exploitable vulnerability. Therefore, adversarially diverting the trajectory can disrupt convergence towards the high quality audio region on the manifold.
@@
Additionally, we target the U-Net’s internal feature representations as a vulnerability for degrading perceptual quality. Prior attribute editing work (Ronneberger, Fischer, and Brox 2015; Oh, Lee, and Lee 2024; Tumanyan et al. 2023) confirms that these hierarchical features within the U-Net are controllable and also encode crucial semantic and acoustic details that govern the coherence and naturalness of the synthesized speech. Therefore, adversarially corrupting these representations directly impair the model’s ability to synthesize natural-sounding speech, offering a complementary strategy for fidelity degradation. This analysis reveals critical vulnerabilities in DMs, providing targeted avenues for adversarial intervention.
4 Methodology
@@
Figure 2: Overview of the proposed framework. Perturbation optimization is guided by gradients from $\mathcal{L}_{total}$, aggregating four targeting two objectives: (1) Identity Obfuscation (via Opposite-Gender Centroid Guidance and Attention Context Divergence) and (2) Perceptual Fidelity Degradation (Score Magnitude Amplification and Noise-Guided Semantic Corruption).
@@
To provide the necessary background, the problem formulation is firstly established in Section 4.1. As shown in Figure 2, the VoiceCloak consists of four sub-modules, which will be introduced separately next.
4.1 Problem Formulation
@@
Assume that malicious users can obtain reference audio $x_{ref}$ of a target speaker. Leveraging open-source voice cloning models, they can synthesize speech which mimics the vocal characteristics of $x_{ref}$, denoted $\mathcal{VC}(x_{src},x_{ref},t)$. Our goal is to proactively safeguard $x_{ref}$ against unauthorized voice cloning. We achieve this by introducing an imperceptible adversarial perturbation $\delta$ to create a protected version $x_{adv}=x_{ref}+\delta$. This perturbation is optimized to disrupt the diffusion synthesis process when conditioned on $x_{adv}$. Formally, we aim to find an optimal $\delta$ that maximizes the dissimilarity between the outputs generated using $x_{ref}$ and $x_{adv}$.
$\arg\max_{\delta}\ \mathcal{D}(\mathcal{VC}(x_{src},x_{ref},t),\ \mathcal{VC}(x_{src},x_{adv},t)),$ (4)
$\text{subject to}\quad\|\delta\|_{\infty}\leq\epsilon,$
@@
where $\mathcal{D}(\cdot)$ measures the output discrepancy, $t$ is the diffusion timestep, and $\epsilon$ is the $l_{\infty}$-norm budget for the perturbation $\delta$. Our approach focuses on designing specific adversarial objectives that implicitly define $\mathcal{D}$ by exploiting intrinsic vulnerabilities within the diffusion mechanism itself.
4.2 Adversarial Identity Obfuscation
Opposite-Gender Embedding Centroid Guidance